OpenStudy (anonymous):

. Calculate the number of moles of excess reactant that will be left-over when 50.0 g of KI react with 50.0 g of Br2: 2KI + Br2  2KBr + I2 a. both reactants are consumed completely b. 0.145 moles c. 0.0117 moles d. 0.162 moles

OpenStudy (anonymous):

i have figured out the 50.0gKI and 50.0gBr2...KI=35.8gKBr and Br2 = 74.5gKBr..just don't know what to do from there.

OpenStudy (anonymous):

\[\large n=\frac{m}{M}\] \[\large n_{KI}=\frac{50}{39.099+126.9}\] \[\large =0.3012066mol\] \[\large n_{Br_{2}}=\frac{50}{2(79.9)}\] \[\large =0.3128911mol\]

OpenStudy (anonymous):

I'm confused. I'm not sure what all of that means.

OpenStudy (anonymous):

according to my answer sheet. the answer is D..0.162moles

OpenStudy (anonymous):

You calculate the number of moles for both reactants.

OpenStudy (anonymous):

You find the limiting reagent. The one with the least number of moles.

OpenStudy (anonymous):

And then you sue your balanced equation to calculate the number of moles required for the other reactant. SO for your given example, the limiting reagent is KI. Therefore using the equation you have their. 2moles of KI= 1 mole of Br2 Therefore, the number of moles for BR is half that of KI.

OpenStudy (anonymous):

You use*

OpenStudy (anonymous):

Yeah that's correct.

OpenStudy (anonymous):

i'm sorry, i'm still confused

OpenStudy (anonymous):

So you halve the number of moles for KI. \[\large n_{KI}=\frac{0.3012066}{2}\] \[\large =0.1506033169mol\] \[\large n_{Br_{2}}-n_{KI}=0.3128911-0.1506033169\] \[\large =0.1623mol\] \[\large =0.162 mol (3-significant-figures)\]

OpenStudy (anonymous):

what does 'n' stand for?

OpenStudy (anonymous):

The limiting reagent which is KI in this case, is the one you want in your equation. n stands for number of moles. Don't you know how to calculate the number of moles for a substance?

OpenStudy (anonymous):

I gave you the equation at the very start. If you're unsure of that, you must revise on that equation, because you're not going to know how to do these questions without it.

OpenStudy (anonymous):

yes i do. what does your lower case 'm' stand for and what does you upper case "M" stand for??

OpenStudy (anonymous):

i can calculate, i am unfamiliar with the abbreviations.

OpenStudy (anonymous):

Those are the universal abbreviations for that equation.

OpenStudy (anonymous):

i'm sorry.. i guess it wasn't reviewed so what does it stand for?

OpenStudy (anonymous):

If you do know the equation, you know the abbreviations. That tells me straight away that you do not know how to calculate the number of moles for a substance.

OpenStudy (anonymous):

It shouldn't be reviewed. It would of been taught to you before your teacher gave you these questions.

OpenStudy (anonymous):

n= no. of moles. m= mass of substance M= Molar mass of substance

OpenStudy (anonymous):

thank you.